# Spike No More: Stabilizing the Pre-training of Large Language Models

### 28 Dec 2023 v1

@@ Abstract
The loss spike often occurs during pre-training of a large language model. The spikes degrade the performance of a large language model, and sometimes ruin the pre-training. Since the pre-training needs a vast computational budget, we should avoid such spikes. To investigate a cause of loss spikes, we focus on gradients of internal layers in this study. Through theoretical analyses, we introduce two causes of the exploding gradients, and provide requirements to prevent the explosion. In addition, we introduce the combination of the initialization method and a simple modification to embeddings as a method to satisfy the requirements. We conduct various experiments to verify our theoretical analyses empirically. Experimental results indicate that the combination is effective in preventing spikes during pre-training.
@@ 1 Introduction
Figure 1: Training loss values of Transformers, whose dimensions and the number of layers are the same as the 1.7 billion parameters configuration in Narayanan et al. (2021). In Vanilla, some spikes occur at the beginning of the training, and its loss value exploded at about 13000 step.
In this research, we provide theoretical analyses focusing on the loss spike issue during LLM pre-training. We identify two primary factors that lead to the issue: (1) the rapid amplification of the norm of the residual branch during forward propagation and (2) the intensification of gradients before and after layer normalizations (LNs) (Ba et al., 2016) in each layer. Our research also demonstrates that the impact of each factor is closely related to the chosen parameter initialization technique. We provide evidence that initializing parameters using a typical methodology (He et al., 2015) results in gradient explosion attributable to the residual branch. Conversely, a popular initialization method in LLM pre-training, which reduces certain parameters to smaller values, promotes the exploding gradients due to LNs although it mitigates the gradient explosion arising from the residual branch. To prevent both exploding gradients, we describe a requirement and introduce a simple modification to satisfy the requirement.
@@ 2.1 Pre-LN Transformer
In this paper, we mainly focus on the neural architecture used in the GPT series (Radford et al., 2018, 2019; Brown et al., 2020). They use the Pre-LN Transformer (Xiong et al., 2020) which is the de facto standard architecture in recent implementations of Transformers because the architecture is more stable than the original Transformer architecture when we stack many layers (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023). Let $x\in\mathbb{R}^{d}$ be an input of a layer of the Transformer, the layer outputs $y$ with the following equations:
@@
where $\mathrm{LN}$ is the layer normalization function. We call the first terms in Equations (1) and (2), i.e., $x$ and $x^{\prime}$, residual branch. In addition, $\mathrm{FFN}$ and $\mathrm{Attention}$ are defined as follows [^1] To simplify equations, we omit bias terms.:
@@
$\mathrm{head}_{i}(x)$ $=\mathrm{softmax}(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}})XW_{Vi},$ (5)
where $\mathcal{F}$ is an activation function, $\mathrm{concat}$ concatenates input vectors, $\mathrm{softmax}$ applies the softmax function to an input vector, and $W_{1}\in\mathbb{R}^{d\times d_{ffn}}$, $W_{2}\in\mathbb{R}^{d_{ffn}\times d}$, $W_{Qi}\in\mathbb{R}^{d\times d_{head}}$, $W_{Ki}\in\mathbb{R}^{d\times d_{head}}$, $W_{Vi}\in\mathbb{R}^{d\times d_{head}}$, and $W_{O}\in\mathbb{R}^{d\times d}$ are parameter matrices. In addition, we pack the sequence of input vectors into a matrix as $X$ to compute the self-attention.
@@ 2.2 Initialization
Because an initialization method has considerable influence on the training dynamics of neural methods, various initialization methods have been proposed for Transformers (Zhang et al., 2019b; Nguyen & Salazar, 2019; Zhang et al., 2019a; Wang et al., 2022). In this study, we investigate the training stability of Transformers in two kinds of initialization method; we call them the plain initialization and scaled initialization. In the plain initialization, we initialize all parameters with a normal distribution $\mathcal{N}(0,\sigma)$. In the scaled initialization, which is the widely used method in the pre-training of LLMs (Shoeybi et al., 2020; Le Scao et al., 2022; Biderman et al., 2023), we scale $W_{2}$ and $W_{O}$ to small values. Concretely, we follow Shoeybi et al. (2020); Le Scao et al. (2022) and initialize $W_{2}$ and $W_{O}$ with a normal distribution $\mathcal{N}(0,\sigma/\sqrt{2N})$ where $N$ is the number of layers [^2] Biderman et al. (2023) also scaled $W_{2}$ and $W_{O}$ to small values in the initialization, but they used the strategy introduced by Wang & Komatsuzaki (2021) instead of $\mathcal{N}(0,\sigma/\sqrt{2N})$. We discuss the effect of this initialization method in Appendix C.. Following Le Scao et al. (2022), we use $\sqrt{2/5d}$ for $\sigma$, that is initially proposed by Nguyen & Salazar (2019).
@@ 3 Analyses on Exploding Gradients
In this Section, we discuss two causes of the exploding gradients and the ways to prevent them. We indicate that the explosion due to the residual branch occurs when we use the plain initialization described in Section 2.2. Moreover, we indicate that the scaled initialization mitigates the explosion due to the residual branch but causes the explosion due to LNs. To prevent both explosions, we describe a requirement and introduce a modification in the scaled initialization.
3.1 Explosion due to Residual Branch
In this section, we indicate that the exploding gradients due to the residual branch occurs when we use the plain initialization. Moreover, we indicate that the scaled initialization prevents the exploding gradients due to the residual branch.
@@ 3.1 Explosion due to Residual Branch
For both initialization methods, we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma=\sqrt{2/5d}\ll 1$, and thus, the standard deviations of embeddings are much smaller than $1$. Then, we investigate the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ in Equations (3) and (4). If these standard deviations are also much smaller than $1$, the standard deviations of each layer output by Equations (1) and (2) remain a small scale. In contrast, for the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$, if either is close to or larger than $1$, the standard deviation of the residual branch rapidly grows as the layer gets deeper due to Equation (6). This leads the exponential growth of the norms of each layer calculated by Equations (1) and (2), and thus, causes exploding gradients during the back-propagation.
@@
$\mathrm{var}(\mathrm{FFN}(x))$ $=4d\ d\ \frac{2}{5d}\frac{2}{5d},$
@@
Thus, the standard deviation is $\sqrt{16/25}=4/5\simeq 1$. Because at least the standard deviation of $\mathrm{FFN}(x)$ is close to $1$, the plain initialization exponentially increases the standard deviation of the residual branch during forward computation, and then causes exploding gradients during the back-propagation.
@@
$\mathrm{var}(\mathrm{FFN}(x))$ $=4d\ d\ \frac{2}{5d}\frac{2}{5d}\frac{1}{2N},$
@@
Thus, the standard deviation is $\sqrt{8/25N}$. Due to $N>10$ in LLMs, $\sqrt{8/25N}$ is much smaller than 1.
We indicated that the scaled initialization mitigates the increase of the residual branch by $\mathrm{FFN}(x)$. We prove that the scaled initialization also mitigates the increase by $\mathrm{Attention}(x)$. Let consider the standard deviation of Equation (4) in the scaled initialization. Before computing, we consider the standard deviation of each self-attention head computed by Equation (5). Because $X$ is the pack of layer normalized vectors, the variance of $\mathrm{softmax}(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}})X$ reaches the maximum value $1$ when the result of $\mathrm{softmax}(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}})$ is a $1$-hot vector [^3] On the other hand, the variance reaches the minimum value when the result of the $\mathrm{softmax}$ is a uniform distribution.. Thus, we obtain the variance of Equation (5) as follows:
$\mathrm{var}(\mathrm{head}_{i}(x))$ $=\mathrm{var}(\mathrm{softmax}(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}})X)\ d\ \mathrm{var}(W_{Vi}),$
$=\mathrm{var}(\mathrm{softmax}(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}})X)\ d\ \frac{2}{5d},$
@@
$\mathrm{var}(\mathrm{Attention}(x))$ $=\mathrm{var}(\mathrm{head}(x))\ d\ \mathrm{var}(W_{O}),$
$=\mathrm{var}(\mathrm{head}(x))\ d\ \frac{2}{5d}\frac{1}{2N},$
$=\mathrm{var}(\mathrm{head}(x))\ \frac{1}{5N},$
@@
Based on these calculations, the standard deviation of $\mathrm{Attention}(x)$ is also much smaller than $1$ when we use the scaled initialization. These theoretical analyses prove that the scaled initialization prevents the rapid growth of the residual branch, and thus, it prevents the exploding gradients due to the residual branch.
Figure 2: Gradient norms of Transformers at the beginning of the LLM pre-training. To focus on only the explosion due to the residual branch, we prevent the explosion due to LNs by multiplying embeddings by $\sqrt{d}$ as described in Section 3.2.
Figure 2 shows the gradient norms of each layer at the beginning of the LLM pre-training. This figure shows that the gradient norms in the plain initialization rapidly grow as the layer gets shallower. This means that the plain initialization faces the exploding gradients. In contrast, the gradient norms in the scaled initialization remain almost constant values. Therefore, the scaled initialization, which is widely used in LLM pre-training, is effective on preventing the exploding gradients.
However, in this figure, we prevent the explosion due to LNs to focus on only the explosion due to the residual branch. If we remove the modification, the gradient norms in the scaled initialization also explode. We describe the explosion due to LNs and the modification to address it the next section.
@@ 3.2 Explosion due to LNs
$||\frac{\delta\mathrm{LN}(x)}{\delta x}||_{2}=\mathcal{O}(\frac{\sqrt{d}}{||x||_{2}}).$ (12)
As described in Section 3.1, we assume $||x||_{2}=\sqrt{d}\ \mathrm{std}(x)$. Thus we can modify the above equation as follows:
$||\frac{\delta\mathrm{LN}(x)}{\delta x}||_{2}=\mathcal{O}(\frac{1}{\mathrm{std}(x)}).$ (13)
The input vectors of LNs are an embedding at the first layer, and each layer output that is the sum of the residual branch and $\mathrm{FFN}(x)$ or $\mathrm{Attention}(x)$ as in Equations (1) and (2). As described, the standard deviations of embeddings are much smaller than $1$ because we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma\ll 1$. In addition, since the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ are much smaller than $1$ in the scaled initialization as described in Section 3.1, LNs certainly increase gradients due to Equation (13) especially in shallow layers [^4] The each layer output grows as the layer gets deeper, because we add $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ the number of layers times. For example, we can remove $N$ in denominators of Equations (9) and (11) at the final layer. Therefore, the standard deviations of each layer output at deep layers are close to $1$ in comparison with ones at shallow layers..
To solve this issue, we need to make the standard deviations of vectors immediately before LNs close to $1$. We have two choices to achieve this requirement. One is increasing the standard deviation of the residual branch, and the other is increasing the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention(x)}$. However, as described in Section 3.1, if we make the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention(x)}$ close to $1$, the explosion due to the residual branch occurs because the norm of the residual branch rapidly grows. Thus, we make the standard deviation of the residual branch close to $1$.
To make the standard deviation of the residual branch close to $1$, we make the standard deviation of each embedding close to $1$. To achieve this, we introduce two kinds of modification: “Scaled Embed” and “Embed LN”. The Scaled Embed is scaling embeddings with an appropriate value [^5] In this study, we use word tying (Press & Wolf, 2017; Inan et al., 2017). In this configuration, we have to initialize embeddings with a small value to use them in the output layer. Otherwise, we can initialize embeddings with the normal distribution $\mathcal{N}(0,\sigma)$ where $\sigma=1$ to solve the explosion due to LNs.. For example, we multiply embeddings by $\sqrt{d}$, which is introduced in the original Transformer paper (Vaswani et al., 2017) [^6] Although the original Transformer paper introduced this operation, recent implementations ignore this., and then the standard deviations of embeddings become $\sqrt{2/5}$. The Embed LN is applying the LN to embeddings. In fact, Le Scao et al. (2022) reported that the Embed LN strategy prevents loss spikes empirically.
Figure 3 shows the gradient norms of each layer when we use the scaled initialization. This figure indicates that the gradient exponentially increases especially in shallow layers in Vanilla that uses the scaled initialization but does not use any modification to prevent the explosion due to LNs. In contrast, our modifications prevent the exploding gradients because they resolve the explosion due to LNs. Therefore, we require the combination of the scaled initialization and modifications described in this section to prevent both explosions due to the residual branch and LNs.

### 02 Feb 2024 v2

@@ Abstract
Loss spikes often occur during pre-training of large language models. The spikes degrade the performance of large language models and sometimes ruin the pre-training. Since the pre-training needs a vast computational budget, we should avoid such spikes. To investigate the cause of loss spikes, we focus on gradients of internal layers. Through theoretical analyses, we reveal two causes of the exploding gradients, and provide requirements to prevent the explosion. In addition, we propose a method to satisfy the requirements by combining the initialization method and a simple modification to embeddings. We conduct various experiments to verify our theoretical analyses empirically. Experimental results indicate that the combination is effective in preventing spikes during pre-training.
@@ 1 Introduction
Figure 1: Training loss values of Transformers, whose dimensions and the number of layers are the same as the 1.7 billion parameters configuration in Narayanan et al. (2021). In Vanilla, some spikes occur at the beginning of the training, and its loss value exploded at about 13000 steps.
In this research, we provide theoretical analyses focusing on the loss spike issue during LLM pre-training. We identify two primary factors that lead to the issue: (1) the rapid amplification of the norm of the shortcut part in the residual connection during forward propagation and (2) the intensification of gradients before and after layer normalizations (LNs) (Ba et al., 2016) in each layer. Our research also demonstrates that the impact of each factor is closely related to the chosen parameter initialization technique. We provide evidence that initializing parameters using a typical methodology (He et al., 2015) results in gradient explosion attributable to the shortcut. Conversely, a popular initialization method in LLM pre-training, which reduces certain parameters to smaller values, promotes the exploding gradients due to LNs although it mitigates the gradient explosion arising from the shortcut. To prevent both exploding gradients, we describe a requirement and introduce a simple modification to satisfy the requirement.
@@ 2.1 Pre-LN Transformer
This paper mainly focuses on the neural architecture used in the GPT series (Radford et al., 2018, 2019; Brown et al., 2020). They use the Pre-LN Transformer (Xiong et al., 2020) which is the de facto standard architecture in recent implementations of Transformers because the architecture is more stable than the original Transformer architecture when we stack many layers (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023). Let $x\in\mathbb{R}^{d}$ be an input of a layer of the Transformer, where $d$ denotes the dimension of the layer. The layer outputs $y$ with the following equations:
@@
where $\mathrm{LN}$ is the layer normalization function. We call the first terms in Equations (1) and (2), i.e., $x$ and $x^{\prime}$, shortcut. In addition, $\mathrm{FFN}$ and $\mathrm{Attention}$ are defined as follows [^1] To simplify equations, we omit bias terms.:
@@
$\mathrm{head}_{i}(x)$ $=\mathrm{softmax}\left(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}}\right)XW_{Vi},$ (5)
where $\mathcal{F}$ is an activation function, $\mathrm{concat}$ concatenates input vectors, $\mathrm{softmax}$ applies the softmax function to an input vector, and $W_{1}\in\mathbb{R}^{d\times d_{ffn}}$, $W_{2}\in\mathbb{R}^{d_{ffn}\times d}$, $W_{Qi}\in\mathbb{R}^{d\times d_{head}}$, $W_{Ki}\in\mathbb{R}^{d\times d_{head}}$, $W_{Vi}\in\mathbb{R}^{d\times d_{head}}$, and $W_{O}\in\mathbb{R}^{d\times d}$ are parameter matrices, and $d_{ffn}$ and $d_{head}$ are the internal dimensions of FFN and multi-head attention sub-layers, respectively. In addition, we pack the sequence of input vectors into a matrix as $X$ to compute the self-attention.
@@ 2.2 Initialization
Because an initialization method has considerable influence on the training dynamics of neural methods, various initialization methods have been proposed for Transformers (Zhang et al., 2019b; Nguyen & Salazar, 2019; Zhang et al., 2019a; Wang et al., 2022). In this study, we investigate the training stability of Transformers in two kinds of initialization methods; we call them plain initialization and scaled initialization. In the plain initialization, we initialize all parameters with a normal distribution $\mathcal{N}(0,\sigma)$. In the scaled initialization, which is the widely used method in the pre-training of LLMs (Shoeybi et al., 2020; Le Scao et al., 2022; Biderman et al., 2023), we scale $W_{2}$ and $W_{O}$ to small values. Concretely, we follow Shoeybi et al. (2020); Le Scao et al. (2022) and initialize $W_{2}$ and $W_{O}$ with a normal distribution $\mathcal{N}(0,\sigma/\sqrt{2N})$ where $N$ is the number of layers [^2] Biderman et al. (2023) also scaled $W_{2}$ and $W_{O}$ to small values in the initialization, but they used the strategy introduced by Wang & Komatsuzaki (2021) instead of $\mathcal{N}(0,\sigma/\sqrt{2N})$. We discuss the effect of this initialization method in Appendix C.. Following Le Scao et al. (2022), we use $\sqrt{2/5d}$ for $\sigma$, that is initially proposed by Nguyen & Salazar (2019).
@@ 3 Analyses on Exploding Gradients
This section discusses two causes of the exploding gradients and the ways to prevent them. We indicate that the explosion due to the shortcut occurs when we use the plain initialization described in Section 2.2. Moreover, we indicate that the scaled initialization mitigates the explosion due to the shortcut but causes the explosion due to LNs. To prevent both explosions, we describe a requirement and introduce a modification in the scaled initialization.
3.1 Explosion due to Shortcut
In this section, we indicate that the exploding gradients due to the shortcut occur when we use the plain initialization. Moreover, we indicate that the scaled initialization prevents the exploding gradients due to the shortcut.
@@ 3.1 Explosion due to Shortcut
For both initialization methods, we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma=\sqrt{2/5d}\ll 1$, and thus, the standard deviations of embeddings are much smaller than $1$. Then, we investigate the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ in Equations (3) and (4). If these standard deviations are also much smaller than $1$, the standard deviations of each layer output by Equations (1) and (2) remain on a small scale. In contrast, for the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$, if either is close to or larger than $1$, the standard deviation of the shortcut rapidly grows as the layer gets deeper due to Equation (6). This leads to the exponential growth of the norms of each layer calculated by Equations (1) and (2), and thus, causes exploding gradients during the back-propagation.
@@
$\mathrm{var}(\mathrm{FFN}(x))$ $=4d\ d\ \frac{2}{5d}\frac{2}{5d}$
@@
Thus, the standard deviation is $\sqrt{16/25}=4/5\simeq 1$. Because at least the standard deviation of $\mathrm{FFN}(x)$ is close to $1$, the plain initialization exponentially increases the standard deviation of the shortcut during forward computation, and then causes exploding gradients during the back-propagation.
@@
$\mathrm{var}(\mathrm{FFN}(x))$ $=4d\ d\ \frac{2}{5d}\frac{2}{5d}\frac{1}{2N}$
@@
Thus, the standard deviation is $\sqrt{8/25N}$. Recall that $N$ represents the number of layers as defined in Section 2.2, $\sqrt{8/25N}$ is much smaller than 1 for typical LLMs since they usually use $N>10$.
We indicated that the scaled initialization mitigates the increase of the shortcut by $\mathrm{FFN}(x)$. We prove that the scaled initialization also mitigates the increase by $\mathrm{Attention}(x)$. Let us consider the standard deviation of Equation (4) in the scaled initialization. Before computing, we consider the standard deviation of each self-attention head computed by Equation (5). To simplify the equation, let $\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}}$ be $Z$. Then, we can transform the right term of Equation (5) to $\mathrm{softmax}(Z)XW_{Vi}$. Because $X$ is the pack of layer-normalized vectors, the variance of $\mathrm{softmax}(Z)X$ reaches the maximum value $1$ when the result of $\mathrm{softmax}(Z)$ is a $1$-hot vector [^3] On the other hand, the variance reaches the minimum value when the result of the $\mathrm{softmax}$ is a uniform distribution.. Thus, we obtain the variance of Equation (5) as follows:
$\mathrm{var}(\mathrm{head}_{i}(x))$ $=\mathrm{var}(\mathrm{softmax}(Z)X)\ d\ \mathrm{var}(W_{Vi})$
$=\mathrm{var}(\mathrm{softmax}(Z)X)\ d\ \frac{2}{5d}$
@@
$\mathrm{var}(\mathrm{Attention}(x))$ $=\mathrm{var}(\mathrm{head}(x))\ d\ \mathrm{var}(W_{O})$
$=\mathrm{var}(\mathrm{head}(x))\ d\ \frac{2}{5d}\frac{1}{2N}$
$=\mathrm{var}(\mathrm{head}(x))\ \frac{1}{5N}$
@@
Based on these calculations, the standard deviation of $\mathrm{Attention}(x)$ is also much smaller than $1$ when we use the scaled initialization. These theoretical analyses prove that the scaled initialization prevents the rapid growth of the shortcut, and thus, it prevents the exploding gradients due to the shortcut.
Figure 2: Gradient norms of Transformers at the beginning of the LLM pre-training. To focus on only the explosion due to the shortcut, we prevent the explosion due to LNs by multiplying embeddings by $\sqrt{d}$ as described in Section 3.2.
Figure 2 shows the gradient norms of each layer at the beginning of the LLM pre-training. This figure shows that the gradient norms in the plain initialization rapidly grow as the layer gets shallower. This means that the plain initialization faces the exploding gradients. In contrast, the gradient norms in the scaled initialization remain at almost constant values. Therefore, the scaled initialization, which is widely used in LLM pre-training, is effective in preventing the exploding gradients.
However, in this figure, we prevent the explosion due to LNs to focus on only the explosion due to the shortcut. If we remove the modification, the gradient norms in the scaled initialization also explode. We describe the explosion due to LNs and the modification to address it in the next section.
@@ 3.2 Explosion due to LNs
$\left\lVert\frac{\delta\mathrm{LN}(x)}{\delta x}\right\rVert_{2}=\mathcal{O}\left(\frac{\sqrt{d}}{||x||_{2}}\right).$ (12)
As described in Section 3.1, we assume $||x||_{2}=\sqrt{d}\ \mathrm{std}(x)$. Thus, we can modify the above equation as follows:
$\left\lVert\frac{\delta\mathrm{LN}(x)}{\delta x}\right\rVert_{2}=\mathcal{O}\left(\frac{1}{\mathrm{std}(x)}\right).$ (13)
The input vectors of LNs are an embedding at the first layer, and each layer output that is the sum of the shortcut and $\mathrm{FFN}(x)$ or $\mathrm{Attention}(x)$ as in Equations (1) and (2). As described, the standard deviations of embeddings are much smaller than $1$ because we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma\ll 1$. In addition, since the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ are much smaller than $1$ in the scaled initialization as described in Section 3.1, LNs certainly increase gradients due to Equation (13) especially in shallow layers [^4] The each layer output grows as the layer gets deeper, because we add $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ the number of layers times. For example, we can remove $N$ in denominators of Equations (9) and (11) at the final layer. Therefore, the standard deviations of each layer output at deep layers are close to $1$ in comparison with ones at shallow layers..
To solve this issue, we need to make the standard deviations of vectors immediately before LNs close to $1$. We have two choices to achieve this requirement. One is increasing the standard deviation of the shortcut, and the other is increasing the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention(x)}$. However, as described in Section 3.1, if we make the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention(x)}$ close to $1$, the explosion due to the shortcut occurs because the norm of the shortcut rapidly grows. Thus, we make the standard deviation of the shortcut close to $1$.
To make the standard deviation of the shortcut close to $1$, we make the standard deviation of each embedding close to $1$. To achieve this, we introduce two kinds of modification: “Scaled Embed” and “Embed LN”. The Scaled Embed is scaling embeddings with an appropriate value [^5] In this study, we use word tying (Press & Wolf, 2017; Inan et al., 2017). In this configuration, we have to initialize embeddings with a small value to use them in the output layer. Otherwise, we can initialize embeddings with the normal distribution $\mathcal{N}(0,\sigma)$ where $\sigma=1$ to solve the explosion due to LNs.. For example, we multiply embeddings by $\sqrt{d}$, which is introduced in the original Transformer paper (Vaswani et al., 2017) [^6] Although the original Transformer paper introduced this operation, recent implementations ignore this., and then the standard deviations of embeddings become $\sqrt{2/5}$. The Embed LN is applying the LN to embeddings. In fact, Le Scao et al. (2022) reported that the Embed LN strategy prevents loss spikes empirically.
Figure 3 shows the gradient norms of each layer when we use the scaled initialization. This figure indicates that the gradient exponentially increases especially in shallow layers in Vanilla that uses the scaled initialization but does not use any modification to prevent the explosion due to LNs. In contrast, our modifications prevent the exploding gradients because they resolve the explosion due to LNs. Therefore, we require the combination of the scaled initialization and modifications described in this section to prevent both explosions due to the shortcut and LNs.